# VideoLifter: Lifting Videos to 3D with Fast Hierarchical Stereo Alignment

### v1

@@ 3.1 Fragment Registration with Learned 3D Priors
Despite this refinement, scale variations may persist within the point clouds of the fragment due to independent inference. To address this problem, intersected 3D points from the key frame are utilized for scale estimation across all image pairs. Specifically, for each image pair $(I_{1},I_{i})$ between $I_{1}$ and $\{I_{i}\}_{i=2}^{k}$, the corresponding 3D point positions (with a total of $P$ points) are retrieved. A one-degree-of-freedom (1-DoF) scale factor is computed between the intersected 3D points in the current pair and those from the key frame:
$\small s_{i}=\arg\min_{s}\sum_{n=1}^{P}\|s\mathbf{p}_{n}^{(i)}-\mathbf{p}_{n}^{(1)}\|^{2}.$ (2)
This scale factor is then applied to the dense 3D points of view $I_{i}$ in the next stage, ensuring local alignment of the point clouds within the fragment and maintaining a consistent scale relative to the key frame. Specifically, the point map of $I_{1}$ is treated as stereo metric depth, since it has already been globally optimized among all key frames, maintaining consistency across fragments.
@@ 3.2 Hierarchical Gaussian Alignment
We then perform dense 3D scene reconstruction using Gaussian splatting, ensuring that the final 3D model is both globally coherent while finely capturing local scene details through hierarchical Gaussian alignment.
Specifically, a set of local 3D Gaussians $\mathcal{G}^{f}=\{G^{f}_{i}\}_{i=1}^{m}$ is initialized based on the results from fragment registration. These local Gaussians are progressively merged into larger sets through a pairwise process. However, naive pairwise merging introduces two key challenges: an excessive number of Gaussians for optimization and inconsistencies between different local Gaussian sets. To overcome these issues, we propose a hierarchical pipeline that iteratively performs local Gaussian optimization, cross-fragment alignment, and visibility masking for pairwise merging, culminating in a globally consistent 3D scene.
Local 3D Gaussian Construction
We initialize a distinct set of Gaussians for each fragment, resulting in a sequence of Gaussian sets $\mathcal{G}^{f}=\{G^{f}_{i}\}_{i=1}^{m}$, where $G^{f}_{i}$ is initialized from fragment $\mathcal{I}^{f}_{i}$.
In the fragment registration stage, we obtain the dense point cloud for the key frame, along with the relative pose and scale for other frames in the fragment. Using this information, we create a locally consistent dense point cloud by applying the same scale across the entire point map generated by MASt3R. To initialize $G^{f}_{i}$, we assign a Gaussian to each point in the pixel-wise point cloud, setting its attributes as follows: color based on the corresponding pixel, center at the 3D point location, opacity adhering to the 3D-GS protocol [18], and scaling such that it projects as a one-pixel radius in the 2D image (achieved by dividing the depth by the focal length). Here we set the Gaussians as isotropic to reduce the degrees of freedom in Gaussian training.
Since the initial camera poses and point cloud positions may contain minor inaccuracies, we further refine them through joint optimization of camera poses and Gaussian parameters. Specifically, for each local Gaussian in $G^{f}_{i}$, we randomly sample frames within the fragment, render the current Gaussians into sampled frame, and backpropagate gradient updates to the Gaussian positions, colors, scales, opacities, and camera poses.
Hierarchical Merging with Cross Fragment Alignment
To achieve a globally consistent 3D reconstruction for the entire video, we integrate the sequence of local 3D Gaussian sets $\mathcal{G}^{f}$, each initialized and optimized independently within disjoint fragments, into a cohesive representation. To mitigate incremental error accumulation and manage the number of Gaussians during merging, we employ a hierarchical merging strategy as illustrated in Fig. 3.
Figure 3: Hierarchical Gaussian Alignment. The process iteratively performs three stages: 1) joint optimization of camera poses and local Gaussians (purple), 2) cross-fragment alignment for new local Gaussian (blue), and 3) visibility masking and pairwise merging of local Gaussians (yellow), until a globally consistent scene reconstruction is achieved.
@@
To merge independently optimized $G^{f}_{1}$ and $G^{f}_{2}$, we first perform cross-fragment alignment, using $G^{f}_{1}$ as the reference coordinate system. Within each fragment, the first frame (key frame) has an identity pose, while other frames are defined by their relative poses to the key frame. Using the initial transformation $T^{f}_{1\rightarrow 2}$, we compute the camera poses for the novel frames covered by $G^{f}_{2}$. By enforcing photometric loss on the next novel view while freezing all parameters in $G^{f}_{1}$, we could further optimize $T^{f}_{1\rightarrow 2}$ into $T^{f*}_{1\rightarrow 2}$. Then, we align the Gaussians in $G^{f}_{2}$ with the coordinate system of $G^{f}_{1}$ by using ${T^{f*}_{1\rightarrow 2}}^{-1}$.
To avoid duplicating Gaussians in regions where $G^{f}_{1}$ already provides adequate scene geometry representation on novel view, we use a visibility mask to determine areas that $G^{f}_{1}$ could faithfully reconstruct:
@@
where $\mathbf{p}$ represents pixel position on image plane, $D(\mathbf{p})$ is rendered depth. $\text{Conf}(\mathbf{p})$ is the rendered confidence, e.g. if a pixel contains information from the current gaussian, and threshold $\beta$ determines the masking criteria.
As in Fig. 4, the visibility mask effectively identifies regions where Gaussians from $G^{f}_{1}$ could provide sufficient depth and confidence. Importantly, the mask remains robust to occlusions, even when covered by large Gaussians. Since our initialization is pixel-wise point clouds, the inverse of the visibility mask can be directly applied to $G^{f}_{2}$ to select Gaussians that complement the missing regions of $G^{f}_{1}$. These selected Gaussians inherit previously optimized parameters, ensuring seamless integration into the unified representation.
After merging pair of local Gaussians, further joint optimization is needed to ensure the merged Gaussian set $G^{2f}_{1}$ meets our consistency objectives. We jointly optimize Gaussian properties and camera poses for $G^{2f}_{1}$. Specifically, we randomly sample frames within $[I_{1},I_{8}]$, render all Gaussians into each frame, and backpropagate gradients to Gaussian positions, colors, scales, opacities, and camera poses.
@@ 4.1 Experimental Setup
Datasets
We carry out comprehensive experiments using various real-world datasets, including Tanks and Temples [19], CO3D-V2 [31]. For Tanks and Temples, following NoPe-NeRF [6] and CF-3DGS [12], we assess both novel view synthesis quality and pose estimation accuracy across 8 scenes, which span both indoor and outdoor environments. In each case, we use 7 out of every 8 frames from the video clips as training data and evaluate the novel view synthesis on the remaining frame except Family. For CO3D-V2, which contains thousands of object-centric videos where the camera orbits the object, recovering camera poses is significantly more challenging due to the complex and large camera motions. We follow the protocol of CF-3DGS [12] to choose the same 10 scenes from different object categories and apply the same procedure to divide the dataset into training and testing sets.
@@
Table 1: Quantitative Evaluations on Tanks and Temples Dataset. Our method achieves superior rendering quality and pose accuracy while requiring minimal training time and no camera parameters. (-) indicates no camera parameters required, GT K indicates known intrinsics, GT K & Pose indicates both known intrinsics and extrinsics. The best results are highlighted in bold.
Metrics
We assess our approach on two key tasks using benchmark datasets [19, 31, 24]: generating novel views and estimating camera poses. For the task of novel view synthesis, we evaluate performance using common metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [41], and Learned Perceptual Image Patch Similarity (LPIPS) [45]. In terms of camera pose estimation, we evaluate Absolute Trajectory Error (ATE) [38] and utilize COLMAP-generated poses from all dense views as our ground-truth for evaluation. While Relative Pose Error (RPE) [38] evaluates the local consistency of relative transformations between consecutive frames, it can be sensitive to discrepancies in intrinsic parameters such as focal length. However, ATE provides a more comprehensive measure of global trajectory accuracy and is better aligned with the goals of our method, which emphasizes globally consistent 3D reconstruction [38]. As such, we prioritize ATE as the primary metric for evaluating the poses of VideoLifter.
Implementation Details
Our implementation is built on the PyTorch platform. During fragment registration, each fragment consists of $k=4$ frames. For depth map prediction, we utilize MASt3R with a resolution of 512 on the longer side. We run 200 iterations for key frame optimization. For hierarchical Gaussian alignment, we initialize each local Gaussian using the number of pixels within the fragment and train it for 200 steps. Camera poses are represented in quaternion format. For pair-wise merging, the transformation matrix from key frame optimization is applied to the camera poses and Gaussian points of the subsequent local Gaussian. First, the camera poses are optimized with a learning rate of 1e-3 for 200 steps. Next, a mask is rendered to identify inadequately reconstructed regions, where new Gaussians are added. This process is repeated iteratively until a globally consistent Gaussian representation is achieved. We uniformly sample $\frac{1}{2}$ and $\frac{1}{4}$ training views on Tanks and Temples and CO3D-V2, respectively. All experiments were conducted on a single Nvidia A6000 GPU to maintain fair comparison.
@@ 4.2 Quantitative Evaluations
Table 2: Quantitative Evaluations on CO3D-V2 Datasets. Our method significantly decreases training time in CO3D-V2. (-) indicates no camera parameters required, GT K indicates known intrinsics.The best results are highlighted in bold.
@@ 4.3 Qualitative Evaluations
# Method Train Time SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
1 MASt3R MVS initialization 35min 0.8582 27.91 0.1768
2 No Decomposition 30min15s 0.7563 23.03 0.3515
3 No Hierarchy 53min12s 0.6969 20.22 0.3876
4 $k$ = 2 35min2s 0.8936 29.77 0.2138
5 $k$ = 8 38min5s 0.8787 27.51 0.2743
6 $\beta$ = 0.5 23min18s 0.6529 18.50 0.4457
7 $\beta$ = 0.99 43min55s 0.8325 29.08 0.2691
8 Ours($k$ = 4, $\beta$ = 0.9) 28min49s 0.8957 30.02 0.1745
@@ 4.4 Ablation Study
We conducted a series of ablation studies to validate the effectiveness of our design choices in VideoLifter, transitioning from an all-at-once optimization approach using MVS as initialization to a structured two-stage pipeline with hierarchical Gaussian alignment. Results are in Tab. 3.
Quality of directly using MASt3R MVS as initialization? We assess the accuracy of camera poses derived from the MVS point map. As illustrated in row 1 and row 8, there exists considerable scope for enhancing the MVS-initialized scene geometry.
Effect of decomposing the reconstruction pipeline into two stages? As in row 2 and row 8, decomposing a monolithic pipeline into two stages could disentangle the problem and simplify the optimization of scene reconstruction and camera pose.
Hierarchical Gaussian alignment design? As in row 3 and row 8, it confirms that the progressive merging of local Gaussian representations significantly reduces drift, achieving more globally consistent reconstructions across extended video sequences.
Effect of different sizes for local fragments? Smaller fragments yielded more precise local alignment but increased computational overhead, while larger fragments reduced processing time at the cost of minor accuracy loss. We selected an optimal fragment size as 4 to balance computational efficiency and alignment accuracy.
Effect of different threshold of rendering confidence? A smaller $\beta$ allows fewer new Gaussians to merge into the current local Gaussian and degrades the performance, while a too-high $\beta$ adds too many Gaussians and makes the training slower and harder. To balance the number of Gaussians and training time, we set $\beta$ = 0.9 by default.

### v2

@@ 3.2 Fragment Registration with Learned 3D Priors
Scale factor: Scale variations may persist within the point clouds of the fragment due to independent inference. To address this, intersected 3D points from the key frame are utilized for scale estimation across all image pairs. Specifically, for each image pair $(I_{1},I_{i})$ between $I_{1}$ and $\{I_{i}\}_{i=2}^{k}$, the corresponding 3D point positions (with a total of $P$ points) are retrieved. A one-degree-of-freedom (1-DoF) scale factor is computed between the intersected 3D points in the current pair and those from the key frame:
$\small s_{i}=\operatorname{median}(\left\{\|\mathbf{p}_{n}^{(1)}\|/\|\mathbf{p}_{n}^{(i)}\|\right\}_{n=1}^{P}),$ (1)
which can be solved analytically (i.e., by taking the median) with no optimization need. This scale factor is applied to the dense pointmaps of $I_{i}$ in the subsequent stage, ensuring that the point clouds within the fragment are locally aligned and maintain a consistent scale relative to the key frame.
By solving a simplified optimization problem rather than naively employing MASt3R global optimization as in InstantSplat, our method achieves a processing time of 2.97 seconds compared to 10.33 seconds—a 3× reduction on computational time within fragment $k$=4.
# Views MASt3R (Global Optimization) VideoLifter (Sec. 3.2)
Time (min) Peak GPU Mem (GB) Time (min) Peak GPU Mem (GB)
32 11 4.75 1.7 4.45
48 33 8.86 2.6 5.43
64 63 14.44 3.5 6.38
128 OOM OOM 7.9 16.9
Table 2: Time and peak GPU memory usage on A6000 for varying view counts: MASt3R global optimization vs. our Sec. 3.2.
Overall Efficiency/Quality Improvement with Our Novel Fragment Registration: In contrast to naively applying MASt3R’s global optimization as InstantSplat [12] to compute point clouds and camera poses for all images, our method significantly enhances efficiency. As in Tab. 2, their global optimization approach runs out of memory beyond 64 views and incurs up to an 18$\times$ longer inference time. Moreover, Tab. 5 demonstrates that replacing our fragment registration with MASt3R initialization yields lower-quality results. This is primarily because it trys to solve a more complex global problem with less accurate pointmaps/poses outputs, whereas our approach eases optimization complexity by providing a more robust initialization for subsequent 3D-GS, delivering better quality.
@@ 3.3 Hierarchical Gaussian Alignment
In this stage, we perform dense 3D scene reconstruction using Gaussian splitting. First, we construct local 3D Gaussians within each fragment, and then merge these local models via hierarchical Gaussian alignment. The key design question is how to construct a globally coherent 3D scene while preserving local scene details without incurring significant alignment errors.
Local 3D Gaussian Construction:
We initialize a set of Gaussians, denoted as $\mathcal{G}^{f}=\{G^{f}_{i}\}_{i=1}^{m}$, where $G^{f}_{i}$ is independent initialized and optimize from fragment $\mathcal{I}^{f}_{i}$.
Guassian initialization: In the local fragment registration step, we obtain the key frame’s dense point cloud, along with the relative poses and scale factors for the other frames, which can be used to obtain entire point map within fragment. To initialize $G^{f}_{i}$, we then assign a Gaussian to each point in the pixel-wise point cloud, setting its attributes as: color based on the corresponding pixel, center at the 3D point location, opacity adhering to the 3D-GS protocol [21], and scaling such that it projects as a one-pixel radius in the 2D image (by dividing the depth by the focal length). We set Gaussians as isotropic to reduce the degrees of freedom in Gaussian training.
Further refinement: The initial camera poses and point cloud positions may contain minor inaccuracies, we further refine them through joint optimization of camera poses and Gaussian parameters. Specifically, for each local Gaussian in $G^{f}_{i}$, we randomly sample frames within the fragment, render the current Gaussians into sampled frame, and backpropagate gradient updates to the Gaussian positions, colors, scales, opacities, and camera poses.
Figure 4: Hierarchical Gaussian Alignment. The process iteratively performs three stages: 1) joint optimization of camera poses and local Gaussians (pink), 2) cross-fragment alignment for new local Gaussian (purple), and 3) visibility masking and pairwise merging of local Gaussians (yellow), until a globally consistent scene reconstruction is achieved.
@@
Hierarchical Gaussian Alignment:
Next, we merge the local fragment-level 3D Gaussian sets $\mathcal{G}^{f}=\{G^{f}_{i}\}_{i=1}^{m}$ to the final consistent 3D scene. Naive pairwise progressive merging poses Challenge ➀: an excessive number of Gaussians for optimization, and Challenge ➁: inconsistencies among the various local Gaussian sets. To avoid these issues, we propose a tree-based hierarchical pipeline (see Fig. 4) that iteratively performs three key processes.
1) Inter-Fragment Alignment with Key Frame Guidance (Fig. 4 purple): To merge two independently optimized fragments (e.g., $G^{f}_{1}$ and $G^{f}_{2}$), we first perform cross-fragment alignment to ensure a faithful merging by using $G^{f}_{1}$ as the reference coordinate system. In each fragment, the key frame, i.e., the first frame, is assigned an identity pose, and the remaining frames are defined by their relative poses to this key frame. In Section 3.2, we enforce consistency between key frames and obtain the initial transformation $T^{f}_{1\rightarrow 2}$. We then use this information as a guide to compute the camera poses for the novel frames covered by $G^{f}_{2}$. By enforcing photometric loss on the next novel view while freezing all parameters in $G^{f}_{1}$, we could further optimize $T^{f}_{1\rightarrow 2}$ into $T^{f*}_{1\rightarrow 2}$. Then, we align the Gaussians in $G^{f}_{2}$ with the coordinate system of $G^{f}_{1}$ by using ${T^{f*}_{1\rightarrow 2}}^{-1}$. As shown in Tab. 5, omitting key frame guidance leads to prolonged optimization and degraded performance.
2) Pair-wise merging with visibility-mask-driven Gaussian pruning (Fig. 4 yellow): To avoid duplicating Gaussians in regions where $G^{f}_{1}$ already provides adequate scene reconstruction, we use a visibility mask to determine areas that $G^{f}_{1}$ could faithfully reconstruct:
@@
where $\mathbf{p}$ represents pixel position on image plane, $D(\mathbf{p})$ is rendered depth. $\text{Conf}(\mathbf{p})=\sum_{i}\alpha_{i}(\mathbf{p})\prod_{j=1}^{i-1}\left(1-\alpha_{j}(\mathbf{p})\right)$ is the rendered confidence, e.g. if a pixel contains information from the current gaussian, and threshold $\beta$ determines the masking criteria.
As in Fig. 5, the visibility mask effectively identifies regions where Gaussians from $G^{f}_{1}$ could provide sufficient depth and confidence. Importantly, the mask remains robust to occlusions, even when covered by large Gaussians. As our initialization is pixel-wise point clouds, the inverse of the visibility mask can be directly applied to $G^{f}_{2}$ to select Gaussians that complement the missing regions of $G^{f}_{1}$. These selected Gaussians inherit previously optimized parameters, ensuring seamless integration into the unified representation.
3) Joint optimization (Fig. 4 pink): After merging pair of local Gaussians, further joint optimization is needed to ensure the merged Gaussian set $G^{2f}_{1}$ meets our consistency objectives. We jointly optimize Gaussian properties and camera poses for $G^{2f}_{1}$. Specifically, we randomly sample frames within $[I_{1},I_{8}]$, render all Gaussians into each frame, and backpropagate gradients to Gaussian positions, colors, scales, opacities, and camera poses.
@@ 4.1 Experimental Setup
Datasets: We carry out comprehensive experiments on various real-world datasets, including Tanks and Temples [22], CO3D-V2 [35]. For Tanks and Temples, following NoPe-NeRF [6] and CF-3DGS [14], we assess both novel view synthesis quality and pose estimation accuracy across 8 scenes, spanning both indoor and outdoor environments. In each case, we use 7 out of every 8 frames from the video clips as training data and evaluate the novel view synthesis on the remaining frame except Family. For CO3D-V2, containing thousands of object-centric videos where the camera orbits the object, recovering camera poses is significantly more challenging due to the complex and large camera motions. We follow the protocol of CF-3DGS [14] to select the same 10 scenes from different object categories and apply the same procedure to divide training and testing sets.
@@
Table 3: Quantitative Evaluations on Tanks and Temples Dataset. Our method achieves superior rendering quality and pose accuracy while requiring minimal training time and no camera parameters. (-) indicates no camera parameters required, GT K indicates known intrinsics, GT K & Pose indicates both known intrinsics and extrinsics. InstantSplat cannot process dense views directly due to OOM (see Tab. 2); thus, we adopt its chunk-by-chunk version, which yields inferior quality on long-sequence videos.
Metrics: We assess our approach on two key tasks: generating novel views and estimating camera poses. For the task of novel view synthesis, we evaluate performance using common metrics such as PSNR, SSIM [45], and LPIPS [49]. In terms of camera pose estimation, we evaluate Absolute Trajectory Error (ATE) [42] and utilize COLMAP-generated poses from all dense views as our ground-truth for evaluation. While Relative Pose Error (RPE) [42] evaluates the local consistency of relative transformations between consecutive frames, it can be sensitive to discrepancies in intrinsic parameters such as focal length. However, ATE provides a more comprehensive measure of global trajectory accuracy and is better aligned with the goals of our method, which emphasizes globally consistent 3D reconstruction [42]. As such, we prioritize ATE as the primary metric for evaluating the poses of VideoLifter.
Implementation Details: Our implementation is built on the PyTorch platform. During fragment registration, each fragment consists of $k=4$ frames. For depth map prediction, we utilize MASt3R with a resolution of 512 on the longer side. We run 200 iterations for key frame optimization. For hierarchical Gaussian alignment, we initialize each local Gaussian using the number of pixels within the fragment and train it for 200 steps. Camera poses are represented in quaternion format. For pair-wise merging, the transformation matrix from key frame optimization is applied to the camera poses and Gaussian points of the subsequent local Gaussian. First, the camera poses are optimized with a learning rate of 1e-3 for 200 steps. Next, a mask is rendered to identify inadequately reconstructed regions, where new Gaussians are added. This process is repeated iteratively until a globally consistent Gaussian representation is achieved. We uniformly sample $\frac{1}{2}$ and $\frac{1}{4}$ training views on Tanks and Temples and CO3D-V2, respectively. All experiments were conducted on a single Nvidia A6000 GPU to maintain fair comparison.
@@ 4.2 Quantitative Evaluations
Table 4: Quantitative Evaluations on CO3D-V2 Datasets. (-) indicates no camera parameters required, GT K indicates known intrinsics.
@@ 4.3 Qualitative Evaluations
Model Train Time SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
Ours($k$ = 4, $\beta$ = 0.9) 28min49s 0.8957 30.02 0.1745
Local: Use MASt3R MVS Init. [12] 35min 0.8582 27.91 0.1768
Global: Hierarchy $\rightarrow$ sequential 53min12s 0.6969 20.22 0.3876
Global: Remove Key Frame Guidance 35min42s 0.7629 24.35 0.3433
$k$ = 2 35min2s 0.8936 29.77 0.2138
$k$ = 4 28min49s 0.8957 30.02 0.1745
$k$ = 8 38min5s 0.8787 27.51 0.2743
$\beta$ = 0.5 23min18s 0.6529 18.50 0.4457
$\beta$ = 0.9 28min49s 0.8957 30.02 0.1745
$\beta$ = 0.99 43min55s 0.8325 29.08 0.2691
@@ 4.4 Ablation Study
Tab. 5 reports the impact of various design choices on training time and reconstruction quality.
Local Fragment Registration. Replacing it with direct MASt3R multi-view stereo initialization increases training time and lowers reconstruction quality, suggesting that outputs directly from MASt3R is less accurate and introduces errors to gaussian optimization, especially in the challenging long video sequences.
Hierarchical Gaussian Alignment. Removing our hierarchical alignment and instead adding local Gaussians sequentially (as in CF-3DGS [14]) prolongs training and degrades performance, underscoring the efficiency and accuracy gains from our hierarchical design.
Key Frame Guidance. Omitting key frame guidance forces additional time for pose optimization without achieving optimal performance, showing the crucial role of key frames in stabilizing and accelerating the merging process.
Fragment Size $k$. A smaller $k$ yields more precise intra-fragment registration but complicates fragment alignment, whereas a larger $k$ reduces joint correspondences within fragment and degrades relative pose estimation, thus degrading the performance.
Confidence Threshold $\beta$. Setting $\beta$ too low allows fewer Gaussians to merge—leading to under-reconstructed areas—while a high $\beta$ merges too many Gaussians, complicating and slowing down training.